OpenStudy (anonymous):

im suppose to solve x+(9/x+1)=5 help please

OpenStudy (anonymous):

your pretty girl

OpenStudy (anonymous):

thank you very much :)

OpenStudy (anonymous):

i wish i could help you but im not rly sure how to do that problem :( SRY

OpenStudy (anonymous):

its ok :/

OpenStudy (anonymous):

\[[x (x + 1)]/(x + 1) + 9/(x + 1)\] then x^2 + x +9 = 5(x +1) which is x^2 +x + 9 = 5x + 5

OpenStudy (radar):

\[x+(9/(x+1))\] Is that the problem

OpenStudy (radar):

If it is x=2 2+(9/(2+1))=5 2+9/3=5 2+3=5 5=5

OpenStudy (radar):

Do you want to see how it is solved?

OpenStudy (anonymous):

yes please!!!